If a machine can fill 2 cartons in 3 seconds how many cartons can it fill in 10 seconds?
Found 2 solutions by TimothyLamb, Abdullahi2 hassan mamman:
Answer by TimothyLamb(4379)

r = c / t
c = r * t
---
r = 2/3 cartons /sec
---
c = r * time
c = (2/3 cartons/sec) * (10 sec)
c = 20/3
---
c = 6 + 2/3 cartons
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by Abdullahi2 hassan mamman(4)

3 seconds fill 2
10/3 * 2 = 6.6 catons